OpenStudy (anonymous):

prove using induction: 1*1!+2*2!...n*n!=(n+1)!-1

OpenStudy (anonymous):

p(k)=1*1!+2*2!....k*k!=(k+1)!-1 p(k+1)=(k+1)!+1+(k+1)!*(k+1)= (k+2)!-1

OpenStudy (anonymous):

but then i get stuck... i dont know what to do. on the left side i can take out (k+1)! but i dont see how that helps me